### v3

@@ 3.1 Tokens Distributed Over Time
Protocol | Airdrop recipients | # of hops to reach an exchanges | |||||
Total | # sent to exchanges | % sent to exchanges | Mean | Median | Min. | Max. | |
ENS | 83.79% | 1.65 | 2 | 1 | 50 | ||
dYdX | 94.83% | 1.77 | 2 | 1 | 17 | ||
1inch | 96.57% | 1.50 | 1 | 1 | 7 | ||
Gemstone | 99.93% | 1 | 1 | 1 | 1 | ||
Arbitrum | 95.18% | ||||||
Uniswap | 96.75% | 2.21 | 2 | 1 | 17 |
Users typically interact with exchanges to convert one token for another and to sell them. To assess how often users receiving an airdrop sell these tokens on exchanges for profit, we analyzed user interactions with exchanges post-airdrop. Table 3 indicates that a significant proportion of airdrop recipients engage with exchanges to trade their tokens. This engagement ranges from a minimum of 83.79% in the case of ENS to a maximum of 99.93% for Gemstone.
Additionally, Table 3 also presents the shortest path from each recipient address of an airdrop to any exchange address in our dataset. Our findings reveal that the transfer of tokens to exchanges generally involves only a few steps, suggesting that airdrop recipients do not make significant efforts to conceal their activities. Gemstone is a notable outlier, with all tokens being sold within a single hop. Surprisingly, most accounts managed to reach an exchange with a relatively low number of intermediary hops (2 hops). This observation underscores the pivotal role that exchanges play within the cryptocurrency ecosystem.
The majority of accounts exchanged their tokens within approximately one million blocks post-airdrop. The number of blocks for Gemstone is significantly higher than others due to a delay introduced by its developer. Given the difference in block generation times between Ethereum (with a new block created every 15 seconds) and ZKsync, we normalized the blocks in terms of days. From Fig. 1(b), within a single day, 66.09% of the 1inch accounts interacted with an exchange. In comparison, ENS had an interaction rate of 55.15%, while dYdX stood at 64.26% followed by Arbitrum and Uniswap with 60.34% and 12.39%, respectively. This rapid exchange behavior does not align with one of the primary objectives of a protocol conducting an airdrop, which is to promote sustained user engagement, as discussed in §2. The prompt exchange of tokens suggests that users might be leaving a protocol soon after receiving their airdrops.
@@ 3.2 Token Distribution Graphs
To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network. Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another. Specifically, the ENS network consists of nodes and edges, the dYdX network comprises nodes and edges, the Gemstone network encompasses nodes and edges, and the 1inch network features nodes and edges. Lastly, Arbitrum accounts for nodes and edges while Uniswap contains nodes and edges.
To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdrop the first few hours after the airdrop and plot their largest component in Fig. 7 in §0.B. We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer [15]. We show that for all protocols but Gemstone, the DEXthat received the majority of transfers (measured in terms of in-degree) from airdrop recipients is Uniswap, followed by SushiSwap.
In the case of Gemstone, all tokens were sent to the address 0x7aa49ad in just a single hop. On the other hand, in dYdX, the airdrop utilized a more diverse set of exchange addresses. Notably, as depicted in Table 3, a portion of the airdrop recipients chose to sell their tokens on an exchange. We see some common exchanges like Uniswap, Wintermute, and SushiSwap.
@@ 4 Measuring the Airdrop Lift
Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially [16]. Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic [16, 25]. In this section, we examine this through the lens of some relevant metrics such as daily transaction count, daily active addresses, median transaction cost, Total Value Locked (TVL), fees paid by users, and stablecoin market cap applied to the Arbitrum airdrop. We rely on data source from Growthepie [29]. For more details on this data refer to §0.A in the appendix.
@@ 4.1 Unique Active Addresses
Others outperformed Arbitrum without an airdrop. While the number of unique addresses on Arbitrum increased post-airdrop (March 22nd 2023 [57]) and remained 50% higher than pre-airdrop levels, other protocols achieved similar growth without an airdrop. For example, per Fig. 2(b), Optimism saw a larger increase in addresses in May 2023, likely due to the Bedrock launch [54]. Similarly, ZKsync Era surpassed Arbitrum’s address count two months after the airdrop, as shown in Fig. 2(a).
Fees may explain the closing gap between Arbitrum and Optimism. Arbitrum has consistently led Optimism in daily active addresses. However, Fig. 8 in §0.C shows this gap narrowing. Before the airdrop, Arbitrum had 2.6 times more active addresses than Optimism, but in the last 50 days, this has decreased to 1.83 times. Optimism’s lower median transaction fees since June may partly explain this trend (Fig. 3(a)), with the average fee gap between the two protocols being even smaller (Fig. 3(b)).
Unique address counts oscillates over time. The number of unique addresses shows oscillatory behavior, as seen in Fig. 2(b), with rapid rises, peaks, and drops. Notably, Optimism and Arbitrum display inverted phases in relative address counts, possibly due to users switching between protocols as fees rise. However, Fig. 3(a) shows no such pattern in median transaction fees, with Arbitrum consistently having lower fees in mid-May 2023. However, unique address counts metric can be gameable. Per our analysis, Arbitrum’s airdrop did not lead to long-lasting user involvement, as the unique address metric can be gamed. Users can create multiple addresses (Sybil attacks) to exploit airdrop limits. This makes the metric unreliable for measuring genuine activity, as large fluctuations may result from such behavior. Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such activities [10, 24, 53, 63]. Thus, alternative, more Sybil-resistant metrics should be considered to gauge authentic user involvement such as graph network analysis [16, 43] combined with machine learning techniques [31, 50, 66].
@@ 4.2 Transaction-related Metrics
The Arbitrum-Optimism gap narrows when considering transactions. Notably, the transaction count gap between Arbitrum and Optimism had nearly closed by the end of July (see Fig. 4(a)). Additionally, Fig. 4(b) shows that Immutable X’s daily transaction count has almost halved since Arbitrum’s airdrop, whereas its unique address count remained relatively stable. This suggests that while address numbers held steady, user engagement with Immutable X has decreased.
Arbitrum’s per-address transaction count declined post-airdrop. To assess how Arbitrum’s airdrop impacted user engagement in other protocols, Fig. 5(a) shows the relative average daily transaction count per unique address. Since the airdrop, Arbitrum’s per-user transaction count has dropped to less than 75% of its pre-airdrop level. However, since transaction counts can be misleading without considering fees, high transaction counts alone may not indicate genuine engagement. In this regard, some protocols require users to perform multiple transactions to qualify for airdrops, leading to inflated activity when fees are low. This aligns with Goodhart’s law [46], which states that “when a measure becomes a target, it ceases to be a good measure”.
Since June, average transaction fees are similar for all protocols. Good metrics should reflect a user’s commitment, and transaction fees can serve as a proxy for this since it measures how much a user is willing to expend for interacting with the protocol. Since June, the average fee per transaction and per unique address have been similar across protocols. Comparing average and median fees (Fig. 3) shows that the median fee may be more useful in understanding shifts in user behavior. Another useful metric is the relative average fees per address compared to the 50 days before Arbitrum’s airdrop, as shown in Fig. 5(b). It shows that Arbitrum’s user engagement was not significantly impacted by the airdrop and generally follows similar patterns to other protocols.
Arbitrum’s total daily fees spiked around the time of the airdrop. Arbitrum’s airdrop did not provide a significant long-term advantage over its competitors in terms of income from transaction fees as shown in Fig. 9 in §0.C. While Arbitrum experienced a peak in fees on the day of the airdrop, this increase was short-lived. Indeed, per Fig. 6(a), in the 50 days before the airdrop, Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
@@ 4.3 Total Value Locked (TVL)
Arbitrum’s airdrop had a lasting impact on its TVL. Among all examined metrics, TVL is the only one that showed a long-lasting improvement following an airdrop: Arbitrum’s TVL increased by over 50% immediately afterwards, and has not dropped significantly since, as shown in Fig. 6(b). It is perhaps surprising, as Arbitrum’s airdrop distribution strategy considered only user activity performed before February 6, 2023 [19].
@@ 6.3 Proactive Oversight and Community Engagement
During the airdrop process, continuous monitoring and analysis of the protocol’s data are crucial to prevent exploitation. A notable example is provided by Linea, whose team identified a vulnerability that allowed users to manipulate their incentive mechanism [41]. This prompt discovery averted cheaters from claiming over one-third of the Non-Fungible Tokens (NFTs)offered as incentives [58].
@@ 7 Related Work
Airdrop Studies Yaish and Livshits [71] present a theoretical model for airdrops that accounts for both honest users and farmers, where the latter have lower airdrop eligibility costs and lower intrinsic utility from using the platform that issues the airdrop. In their analysis, they show that airdrops in which the issuer expends a non-zero fixed cost for each recipient, the threat posed by farmer sybil attacks lead to unbounded issuance costs. However, they show losses from farmers can be bounded by setting in advance the total quantity of airdrop tokens and splitting them equally among recipients. Moreover, by designing the airdrop mechanism correctly, farmers can be harnessed to foster network effects, thereby attracting honest users who may otherwise would prefer competing platforms.
@@
Technical Aspects of Airdrops Frowis et al. [26] identify the operational challenges and costs associated with large-scale airdrops on Ethereum. They suggest that while cost savings of up to 50% are possible with specific smart contract optimizations, pull-based approaches can shift the cost burden to recipients. Overall, the total cost remains proportional to the number of recipients.
@@ 8 Conclusion
This study identified common issues with airdrops and proposed guidelines to improve their effectiveness. Our analysis of six major airdrops revealed a consistent pattern of recipients quickly selling their tokens—36.62% of ENS, 35.45% of dYdX, and 54.05% of 1inch tokens were traded shortly after distribution, with a median of just two transactions. This suggests airdrops fail to sustain long-term engagement or attract valuable contributors.
@@
For Arbitrum, we observed a spike in daily fees during the airdrop, followed by a decline in transactions per address. Other protocols outperformed Arbitrum without an airdrop, and since June 2023, transaction fees have been similar across protocols, indicating airdrops are not the main driver of user growth.
Lastly, we discussed challenges like airdrop farming, governance token distribution, and insider trading, offering insights for future airdrop strategies.
Acknowledgments
This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.
Appendix 0.A Reproducibility
To ensure full reproducibility of our empirical analysis, we will openly share both our dataset and scripts. The primary data source used in §3 consists of blockchain data extracted from Ethereum, Arbitrum, and ZKsync Era archive nodes. We collected all relevant blocks, transactions, and events tied to the contracts of the six airdrops examined.
For §4, data was sourced from Growthepie.xyz [29] corresponding to the year 2023, which offers various metrics, including daily transaction counts, daily active addresses, median transaction costs, TVL, user fees, and stablecoin market capitalization. These metrics are available for major blockchains such as ZKsync Era, OP Mainnet, Arbitrum, Immutable X, and Polygon zkEVM. To enhance clarity, all metrics were smoothed using a 7-day moving average.
Appendix 0.B Graph Network of the Token Distribution
We construct a graph network (V, E) of the token transfer network for the six airdrops analyzed in this paper and present them in Fig. 7. Each node (V) represents an address, and an edge (E) is created whenever tokens are transferred from one address to another.
For the ENS network, there are nodes and edges. The dYdX network consists of nodes and edges. The Gemstone network contains nodes and edges. The 1inch network includes nodes and edges. The Arbitrum network has nodes and edges, while the Uniswap network contains nodes and edges.
Appendix 0.C More Insights Into Airdrop Metrics

### v4

@@ 4.1.3 Exchange Interaction and Transfer Hops
| Protocol | Airdrop recipients | # of interactions with exchange | # of hops to reach an exchange | ||||||||
| Total | # sent to exchanges | % sent to exchanges | Mean | Median | Min. | Max. | Mean | Median | Min. | Max. | |
| 1inch | 60.24% | 2.64 | 1 | 1 | 2695 | 1.42 | 1 | 1 | 7 | ||
| Arbitrum | 40.36% | 4.11 | 1 | 2 | 1.69 | 2 | 1 | 31 | |||
| Arkham | 15.63% | 1.05 | 1 | 1 | 14 | 2.76 | 2 | 1 | 79 | ||
| DYDX | 22.48% | 2.39 | 1 | 1 | 196 | 1.76 | 2 | 1 | 17 | ||
| ENS | 52.24% | 1.78 | 1 | 1 | 107 | 1.49 | 1 | 1 | 50 | ||
| Lido | 76.55% | 4.21 | 1 | 2 | 132 | 1.17 | 1 | 1 | 4 | ||
| Optimism | 61.12% | 4.71 | 1 | 1 | 1.47 | 1 | 1 | 32 | |||
| TornadoCash | 53.91% | 2 | 1 | 1 | 53 | 1.55 | 1 | 1 | 5 | ||
| Uniswap | 53.77% | 3.10 | 1 | 1 | 1.56 | 1 | 1 | 16 | |||
| Protocol | Airdrop recipients | # of days to transfer to an exchange | ||||||
| Total | # sent to exchanges | % sent to exchanges | Mean | Std. | Min. | Median. | Max. | |
| 1inch | 60.24% | 24.96 | 107.15 | 0 | 0.07 | 1352.73 | ||
| Arbitrum | 40.36% | 61.46 | 107.00 | 0 | 5.13 | 418.37 | ||
| Arkham | 15.63% | 11.37 | 47.71 | 0 | 0.01 | 456.76 | ||
| DYDX | 22.48% | 37.09 | 119.32 | 0 | 0.04 | 1091.45 | ||
| ENS | 52.24% | 69.36 | 172.97 | 0 | 1.04 | 1073.65 | ||
| Lido | 76.55% | 55.32 | 167.47 | 0 | 1.08 | 1099.55 | ||
| Optimism | 61.12% | 70.24 | 137.44 | 0 | 0.30 | 662.77 | ||
| TornadoCash | 53.91% | 46.06 | 114.31 | 0 | 2.86 | 1248.52 | ||
| Uniswap | 53.77% | 32.27 | 119.33 | 0 | 0.11 | 1490.03 | ||
4.1.4 Recipient Engagement and Token Utilization
To assess long-term engagement, we studied token transfer patterns post-airdrop using Table 6. We found that while over 60% of airdrop recipients across all protocols performed at least one transfer, most did not continue interacting beyond the initial transaction.
For example, the median number of token-related transfers per recipient was one across most protocols, with only Lido (median = 2) exceeding this marginally. Likewise, the time span between a recipient’s first and last transfer (a proxy for continued engagement) was short: the median was 0 days for most protocols. Only Lido exhibited a higher median of 42.62 days, reflecting longer engagement likely due to staking or yield incentives.
These findings are consistent with the activity shown in Figure 2 and Figure 4. Specifically, recipients’ token transfer activity peaked within days of the airdrop and dropped sharply thereafter, indicating that tokens were often liquidated in a one-time event rather than incorporated into ongoing protocol usage.
| Protocol | Total of accounts | # of transfers per address | Days between first and last token transfers | ||||||||
| Mean | Std. | Min. | Median | Max. | Mean | Std. | Min. | Median | Max. | ||
| 1inch | (97.61%) | 2.55 | 17.54 | 1 | 1 | 62.76 | 190.18 | 0 | 0 | ||
| Arbitrum | (97.82%) | 4.30 | 89.45 | 1 | 1 | 86.65 | 136 | 0 | 0 | 418.43 | |
| Arkham | (98.94%) | 1.06 | 0.42 | 1 | 1 | 58 | 2.93 | 27.84 | 0 | 0 | 458.57 |
| DYDX | (97.45%) | 2.20 | 3.52 | 1 | 1 | 214 | 84.38 | 198.32 | 0 | 0 | |
| ENS | (90.94%) | 1.70 | 1.83 | 1 | 1 | 117 | 57.79 | 175.26 | 0 | 0 | |
| Lido | (92.9%) | 4.33 | 8.90 | 1 | 2 | 147 | 231.17 | 348.81 | 0 | 42.62 | |
| Optimism | (96.02%) | 4.83 | 51.79 | 1 | 1 | 163.63 | 222.20 | 0 | 0 | 663.21 | |
| Tornado | (91.64%) | 1.85 | 5.59 | 1 | 1 | 327 | 20.17 | 92.57 | 0 | 0 | |
| Uniswap | (98.44%) | 3.73 | 228.29 | 1 | 1 | 64.93 | 195.98 | 0 | 0 | ||
4.1.5 Top Airdrop Earners
In this section, we present the top-1 airdrop recipient for each of the nine protocols analyzed. Airdropped tokens may be distributed to various types of accounts, including Externally Owned Accounts (EOAs), smart contracts, and multi-signature (multisig) wallets. Some of these recipients might be directly affiliated with the protocol—such as treasury wallets or governance-related accounts—while others may represent independent users or airdrop farming operations.
Multisig wallets, which require multiple signatures to authorize transactions, are often used by protocol teams for treasury or operational purposes [59]. Consequently, top recipients frequently include protocol-controlled accounts, as discussed in Table 3.
Table 7 lists the top airdrop recipients for each protocol, along with the number and percentage of tokens received, both in relation to the airdrop and the total token supply. Wallet labels were assigned using the methodology described in §3. We highlight two representative examples. First, the wallet labeled analytico.eth received % of the total tokens distributed by the 1inch airdrop. According to Etherscan, this address holds approximately $7.5 million USD in tokens as of May 2025 [21] and has participated in at least 12 other airdrops [16]. These patterns suggest that it is likely part of a coordinated airdrop farming operation. However, the identity behind this wallet remains unknown, and it is unclear whether it is operated by an individual or a group. In contrast, % of Tornado Cash airdropped tokens were distributed to the Tornado Cash Governance, a common practice that aligns with protocols allocating tokens to support long-term ecosystem development. Although some of these wallets received a significant share of the distributed tokens, their allocations remain relatively small when measured against the total token supply, underscoring the limited concentration of airdrops at the protocol level.
| Protocol | Wallet address | Label | # of tokens | % of dist. tokens | % of total supply |
| 1inch | 0xa0f79b73 | analytico.eth | |||
| Arbitrum | 0x79f02b3f | “luckboxcolonel” on OpenSea | |||
| Arkham | 0xf0df8d2f | Laura Kornelija | |||
| DYDX | 0xa6152d92 | — | |||
| ENS | 0x09049859 | ENS wallet | |||
| Lido | 0x97967632 | — | |||
| Optimism | 0xba74ae07 | tarrence.eth | |||
| Tornado Cash | 0x5efda1ce | Tornado Cash Governance | |||
| Uniswap | 0xdd27c3b3 | Sybil Delegate [76] |
@@ 4.2 Token Distribution Graphs
To better understand how airdropped tokens propagate across users, we construct and analyze token transfer networks for each protocol. We define a token distribution graph as a multidirected graph , where each node represents a unique address, and each directed edge denotes a transfer of tokens from address to address after the airdrop.
These networks vary considerably in scale. The graph for 1inch consists of nodes and edges, while Arbitrum’s network includes nodes and edges. Arkham’s network contains nodes and edges, and DYDX’s has nodes and edges. The ENS network comprises nodes and edges, Lido’s has nodes and edges, and Optimism’s includes nodes and edges. Tornado Cash’s network has nodes and edges, while Uniswap’s network includes nodes and edges.
Since our primary goal is to analyze where airdrop recipients sent their tokens in their first post-airdrop transfer, we restrict our analysis to a 1-hop subgraph that captures only the initial outbound transfers from recipient addresses.
Definition 4.2 (1-hop subgraph).
Let denote the set of addresses that received airdropped tokens directly from the protocol within the initial distribution window. The 1-hop subgraph is defined as:
We visualize the largest weakly connected component of each using the Gephi tool, as shown in Figure 5. To enrich the analysis, we manually annotate high in-degree nodes belonging to centralized and decentralized exchanges with labels from our dataset (§3). Across all protocols, we find that the most frequent recipient of airdropped tokens is Uniswap, followed by SushiSwap. Among centralized exchanges, Binance and Huobi appear as the leading recipients.
@@ 5 Measuring the Airdrop Lift
Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially [23]. Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic [23, 33]. In this section, we examine this through the lens of some relevant metrics such as daily transaction count, daily active addresses, median transaction cost, TVL, fees paid by users, and stablecoin market cap applied to the Arbitrum airdrop. We rely on data sourced from Growthepie [37]. For more details on this data refer to §3.
@@ 5.1 Unique Active Addresses
Others outperformed Arbitrum without an airdrop. While the number of unique addresses on Arbitrum increased post-airdrop (March 22nd 2023 [71]) and remained 50% higher than pre-airdrop levels, other protocols achieved similar growth without an airdrop. For example, per Figure 6(b), Optimism saw a larger increase in addresses in May 2023, likely due to the Bedrock launch [65]. Similarly, ZKsync Era surpassed Arbitrum’s address count two months after the airdrop, as shown in Figure 6(a).
Fees may explain the closing gap between Arbitrum and Optimism. Arbitrum has consistently led Optimism in daily active addresses. However, Figure 12 in §B shows this gap narrowing. Before the airdrop, Arbitrum had 2.6 times more active addresses than Optimism, but in the last 50 days, this has decreased to 1.83 times. Optimism’s lower median transaction fees since June may partly explain this trend (Figure 7(a)), with the average fee gap between the two protocols being even smaller (Figure 7(b)).
Unique address counts oscillates over time. The number of unique addresses shows oscillatory behavior, as seen in Figure 6(b), with rapid rises, peaks, and drops. Notably, Optimism and Arbitrum display inverted phases in relative address counts, possibly due to users switching between protocols as fees rise. However, Figure 7(a) shows no such pattern in median transaction fees, with Arbitrum consistently having lower fees in mid-May 2023. However, unique address counts metric can be gameable. Per our analysis, Arbitrum’s airdrop did not lead to long-lasting user involvement, as the unique address metric can be gamed. Users can create multiple addresses (Sybil attacks) to exploit airdrop limits. This makes the metric unreliable for measuring genuine activity, as large fluctuations may result from such behavior. Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such activities [11, 15, 32, 77]. Thus, alternative, more Sybil-resistant metrics should be considered to gauge authentic user involvement, such as graph network analysis [23, 54] combined with machine learning techniques [40, 62, 80].
@@ 5.2 Transaction-related Metrics
The Arbitrum-Optimism gap narrows when considering transactions. Notably, the transaction count gap between Arbitrum and Optimism had nearly closed by the end of July (see Figure 8(a)). Additionally, Figure 8(b) shows that Immutable X’s daily transaction count has almost halved since Arbitrum’s airdrop, whereas its unique address count remained relatively stable. This suggests that while address numbers held steady, user engagement with Immutable X has decreased.
Arbitrum’s per-address transaction count declined post-airdrop. To assess how Arbitrum’s airdrop impacted user engagement in other protocols, Figure 9(a) shows the relative average daily transaction count per unique address. Since the airdrop, Arbitrum’s per-user transaction count has dropped to less than 75% of its pre-airdrop level. However, since transaction counts can be misleading without considering fees, high transaction counts alone may not indicate genuine engagement. In this regard, some protocols require users to perform multiple transactions to qualify for airdrops, leading to inflated activity when fees are low. This aligns with Goodhart’s law [57], which states that “when a measure becomes a target, it ceases to be a good measure”.
Since June, average transaction fees are similar for all protocols. Good metrics should reflect a user’s commitment, and transaction fees can serve as a proxy for this since it measures how much a user is willing to expend for interacting with the protocol. Since June, the average fee per transaction and per unique address have been similar across protocols. Comparing average and median fees (Figure 7) shows that the median fee may be more useful in understanding shifts in user behavior. Another useful metric is the relative average fees per address compared to the 50 days before Arbitrum’s airdrop, as shown in Figure 9(b). It shows that Arbitrum’s user engagement was not significantly impacted by the airdrop and generally follows similar patterns to other protocols.
Arbitrum’s total daily fees spiked around the time of the airdrop. Arbitrum’s airdrop did not provide a significant long-term advantage over its competitors in terms of income from transaction fees as shown in Figure 13 in §B. While Arbitrum experienced a peak in fees on the day of the airdrop, this increase was short-lived. Indeed, per Figure 10(a), in the 50 days before the airdrop, Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
@@ 5.3 Total Value Locked (TVL)
Arbitrum’s airdrop had a lasting impact on its TVL. Among all examined metrics, TVL is the only one that showed a long-lasting improvement following an airdrop: Arbitrum’s TVL increased by over 50% immediately afterwards, and has not dropped significantly since, as shown in Figure 10(b). It is perhaps surprising, as Arbitrum’s airdrop distribution strategy considered only user activity performed before February 6, 2023 [26].
@@ 7.3 Proactive Oversight and Community Engagement
During the airdrop process, continuous monitoring and analysis of the protocol’s data are crucial to prevent exploitation. A notable example is provided by Linea, whose team identified a vulnerability that allowed users to manipulate their incentive mechanism [52]. This prompt discovery averted cheaters from claiming over one-third of the NFTsoffered as incentives [72].
@@ 8 Related Work
Airdrop Studies. Yaish and Livshits [85] present a theoretical model for airdrops that accounts for both honest users and farmers, where the latter have lower airdrop eligibility costs and lower intrinsic utility from using the platform that issues the airdrop. In their analysis, they show that airdrops in which the issuer expends a non-zero fixed cost for each recipient, the threat posed by farmer sybil attacks lead to unbounded issuance costs. However, they show losses from farmers can be bounded by setting in advance the total quantity of airdrop tokens and splitting them equally among recipients. Moreover, by designing the airdrop mechanism correctly, farmers can be harnessed to foster network effects, thereby attracting honest users who may otherwise would prefer competing platforms.
@@
Technical Aspects of Airdrops. Frowis et al. [34] identify the operational challenges and costs associated with large-scale airdrops on Ethereum. They suggest that while cost savings of up to 50% are possible with specific smart contract optimizations, pull-based approaches can shift the cost burden to recipients. Overall, the total cost remains proportional to the number of recipients.
@@ 9 Conclusion
This study identified key challenges in the implementation of airdrops and proposed strategies to enhance their effectiveness. By analyzing nine major airdrops, we observed a recurring pattern where recipients rapidly sold their tokens—with 65.75% of Lido, 58.67% of 1inch, and 48.21% of Optimism tokens traded shortly after distribution, within just 1-hop. This pattern suggests that current airdrop mechanisms may struggle to maintain long-term engagement and attract valuable contributors.
@@
For Arbitrum, we identified a temporary surge in daily fees during the airdrop, followed by a decline in transactions per address. Notably, other protocols without airdrops performed similarly in transaction volume, and since June 2023, transaction fees have stabilized across all observed protocols. This finding underscores that airdrops alone may not be sufficient to drive sustainable user growth.
We also highlighted several key challenges, including airdrop farming, governance token distribution, and insider trading. Addressing these issues can provide critical insights for designing more effective airdrop strategies. To that end, we propose the following guidelines:
Implement Alternative Incentives. Protocols should consider offering fee discounts for future interactions within the protocol to foster sustained user engagement. Multi-round airdrops can also enable protocols to refine strategies and address previously identified issues. Furthermore, developing more nuanced metrics for airdrop eligibility can mitigate airdrop farming by targeting genuine users rather than opportunistic actors. However, insider trading risks may persist under such frameworks.
Target Reputable Entities. Protocols should prioritize distributing tokens to established entities, such as reputable projects, research groups, tech communities, and students. Funding these entities or allocating airdrops to DAOs developing relevant dApps can foster long-term involvement and align incentives with the protocol’s mission.
Proactive Oversight and Community Engagement. Conducting thorough analyses of past airdrops is key, even those on different chains, to identify potential pitfalls and refine strategies. Maintaining open communication channels with the community can further align protocol developments with user interests.
Align Rewards with Costs. Ensuring that rewards scale proportionately with associated costs can help deter airdrop farmers from exploiting the system. Aligning incentives can also foster a more equitable distribution of tokens to genuinely engaged participants.
In conclusion, these guidelines provide actionable strategies for blockchain protocols, researchers, and practitioners seeking to design more effective airdrop campaigns. Additionally, we release the dataset and scripts used in this study to encourage further research and reproducibility in the field [60]. Furthermore, we believe that exploring the efficacy of “lockdrops,” where tokens are distributed but remain locked for a period, can potentially mitigate rapid sell-offs and promote long-term engagement. We leave this to future work.
Appendix A Gemstone: A Case of Sybil Farming on ZKsync Era
Between July 23 and July 31, 2023, a Sybil farming campaign was carried out on ZKsync Era under the pretense of a token airdrop, known as Gemstone [38, 53]. Unlike legitimate airdrops that aim to incentivize long-term user engagement, Gemstone’s campaign was explicitly designed to exploit anticipated eligibility criteria for a future ZKsync airdrop. The orchestrator employed more than Sybil accounts and deployed a closed-source DEXto simulate user activity. Although marketed as a token airdrop, the Gemstone token had no intrinsic value and remains worthless.
Gemstone stands out due to the scale and structure of its distribution. Approximately 99.53% of the token supply was distributed during the short campaign period—an outlier compared to all other airdrops we examined (as discussed in §4). The airdrop also exhibited an unusually high median number of tokens per recipient, reflecting a heavily concentrated distribution pattern. Notably, per Figure 11, 95% of the distributed tokens were routed through fake exchanges created by the farmer, with nearly all ultimately consolidated into a single address (0x7aa49ad) via a single transaction hop. This rapid and centralized token flow indicates highly coordinated activity, consistent with a Sybil attack rather than legitimate user participation.
Appendix B More Insights Into Airdrop Metrics